OpenStudy (anonymous):

Let f(x) = x + 2 and g(x) = x2 Find ( g o f)(-5) A. -3 B. -10 C. 49 D. 9

OpenStudy (anonymous):

@Catlover5925

myininaya (myininaya):

first we deal with the inside function what is f(-5)?

myininaya (myininaya):

then we will do g(of that result)

OpenStudy (anonymous):

sorry im new to this my teach kinf of threw it on me and was like learn it and didnt really show me how

myininaya (myininaya):

but can you find f(-5) or you don't understand that either?

OpenStudy (anonymous):

im sorry i really dont

myininaya (myininaya):

well f(x)=x+2

myininaya (myininaya):

and if we want to find f(-5) then we just need to replace the x's in that f function right there with -5

myininaya (myininaya):

f(-5)=-5+2=?

OpenStudy (anonymous):

-3

myininaya (myininaya):

yes so f(-5)=-3

myininaya (myininaya):

\[\[(g ^o f)(-5)=g(f(-5)) \\ \] \]

myininaya (myininaya):

so we found the inside f(-5)=-3 so now we need to take care of the outside

myininaya (myininaya):

g(-3)=?

myininaya (myininaya):

used the function that is called g above

OpenStudy (anonymous):

um is it -1

myininaya (myininaya):

g(x)=x^2

myininaya (myininaya):

so g(-3)=?

myininaya (myininaya):

x^2 means x*x

myininaya (myininaya):

so (-3)^2 means (-3)*(-3)

myininaya (myininaya):

so that couldn't be -1

OpenStudy (anonymous):

it would be 9

myininaya (myininaya):

yea